A block of mass m is hanging from the roof by a spring of stiffness k. Now a uniform electric field E is switched on. The charge on the block is Q.


 


1. Max. elongation, during suseguent motions of the block, in the spring??


 


2. Amplitude of SHM performed by the spring-mass system is??

Grade:12

1 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear student,

F=1/2Kx2=qE/m

From here calculate k.

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Sagar Singh

B.Tech, IIT Delhi